Can someone help me solve this problem?
I need to solve the following word problem. Be sure to label the unknowns and to show the equation you use for the solution:
Yuri has a board that is 98 in. long. He wishes to cut the board into two pieces so that one piece will be 10 in. longer than the other. What should the length of each piece be?
Thanks, Ashley

Answer by checkley71(8403)

98=X+(X+10) OR 98=X+X+10 OR 98=2X+10 OR 2X=88 OR X=44 & X+10=54
PROOF 44+54=98 OR 98=98